Given utility function U= X0.5Y0.5 where Px= 12 Birr, Bir, Py=4 Birr and the income of
the consumer is, M= 240 Birr.
A. Find the utility maximizing combinations of X and Y.
D. Calculate marginal rate of substitution of X for Y (MRSX.Y) at equilibrium and inierpiet
your result.
1
Expert's answer
2020-02-10T08:53:47-0500

Solution:

A)


"MU_x=0.5 \frac {y^{0.5}}{x^{0.5}}"


"MU_y=0.5 \frac {x^{0.5}}{y^{0.5}}"

"{\frac{MU_x}{p_x}}=\frac {MU_y}{p_y}"

"x \times p_x+ y \times p_y=M"


"y=3x"

"12 \times x+4 \times y=240"

"x=10, y=30."

D)




"\frac {\partial U}{\partial x \partial y} =\frac {0.25} {(xy)^{0.5}}"


"MRS x.y = \frac {\partial U} {\partial x \partial y} = \frac {0.25}{(10 \times 30)^{0.5}}=0.015"